OpenStudy (lilday12):

Factor the trinomial below. Enter each factor as a polynomial in descending order. x2 + 2x - 24

OpenStudy (lilday12):

x+1 x+12

OpenStudy (lilday12):

x+22

Directrix (directrix):

Find two numbers that multiply to -24 and add to 2 --> -6 and 4 x^2 + 2x - 24 = x^2 + 6x - 4x -24 = x(x + 6) - 4(x + 6). Factor out the common factor of (x + 6) to get: (x + 6) ( x - 4) @lilday12

Directrix (directrix):

Questions?

OpenStudy (lilday12):

Factor the trinomial below. Enter each factor as a polynomial in descending order. x2 + 4x + 4

Directrix (directrix):

x2 + 4x + 4 Two numbers that multiply to 4 and add to 4 are 2 and 2. x2 + 4x + 4 = x^2 + 2x + 2x + 4 = x(x+2) + 2(x + 2). (x + 2) is the common binomial factor. = (x + 2) (x + 2) which is sometimes written as (x + 2)^2

OpenStudy (lilday12):

Factor the trinomial below. Enter each factor as a polynomial in descending order. x2 + 6x + 9

OpenStudy (lilday12):

Would it be x+3 x+3